A cross is made between a pea plant with round, yellow seeds

(WwGg) and one with wrinkled, green seeds (wwgg). The

progeny that resulted had the following traits:

round, green: 82

round, yellow: 18

wrinkled, green: 10

wrinkled, yellow: 90

1. Which alleles were on which chromosome in the doubly

heterozygous parent?

2. What is the distance in cM between these two genes?
Answer #1

Parents are round, yellow and wrinkled green

= WwGg. X. wwgg

Progeny are:

round, green: 82 Wwgg

round, yellow: 18 WwGg

wrinkled, green: 10 wwgg

wrinkled, yellow: 90 wwGg

The genes we got are in trans combination because we have got less parental and more of recombinants.

Parental types : round yellow 18

and wrinkled green 10.

Recombinants: round green 82

Wrinkled yellow 90

Total plants: 200

So the gene round and green are on one chromosome and wrinkled yellow are on another chromosome.

That means the parental one we got above are actually due to recombination in trans combination.

Recombination frequency = 18+10/200= 28/200

= 14%.